# Diffusion Autoencoders for Few-shot Image Generation in Hyperbolic Space

### v1

@@ 4.4 Few-Shot Image Generation
As described in Sec. 3, our approach enables few-shot image generation by varying the stochastic subcode $z_{T}$ or shifting the latent code in a randomly chosen semantic direction within the category cluster. We set $r_{\mathbb{D}}=5.5$ and an encoding strength of $0.95$ (i.e., encoding $5\%$ of the information) to achieve this, as illustrated in Fig. 3. We conduct three experiments demonstrating that HypDAE achieves promising few-shot generation, with additional examples in the SM.
Hyp Radius 6.2 6.0 5.5 5.0 4.5
FID($\downarrow$) 15.18 14.67 14.31 14.56 14.71
LPIPS($\uparrow$) 0.7035 0.7278 0.7415 0.7643 0.7935
@@
Quantitative Comparison with State-of-the-Art. To assess fidelity and diversity, we compute FID [26] and LPIPS [74] following the one-shot settings in [41]. Results in Tab. 1 show significant improvements in FID and LPIPS scores on all four datasets except for the FID score on vggfaces due to the low resolution, affirming the advantages of our diffusion-based approach over previous GAN-based methods.
Qualitative Evaluation. We compare HypDAE qualitatively with WaveGAN [71] and HAE [41]. As seen in Fig. 8, HypDAE generates highly diverse images with detailed fidelity, such as fine feather details, which other methods lack or struggle to maintain. A user study further validates our method by rating quality, fidelity, and diversity. Tab. 3 shows HypDAE surpasses all baselines, attributed to its strong retention of identity features and fine-grained detail variation. Additional details are provided in SM.
Text-Guided Image Editing. HypDAE also supports text-guided few-shot generation, a capability lacking in prior methods. Results in Fig. 12 show that the hyperbolic delta mapper achieves accurate, disentangled editing without changing the category. By combining stochastic subcode variation and hyperbolic semantic shifts, HypDAE generates category-consistent images with specific features. More details are available in SM.
@@ 5 Conclusion
In this work, we present HypDAE, the first diffusion-based method for editing hierarchical attributes in hyperbolic space. By learning the semantic hierarchy of images, our diffusion autoencoder enables continuous and flexible editing of hierarchical features. Experimental results show that HypDAE significantly outperforms existing GAN-based approaches, excelling in both random and customized few-shot generation. Furthermore, it facilitates hierarchical image generation with varying semantic similarity, advancing the state of the art in few-shot image generation. We hope this work inspires future research in this direction.
@@
Sec. 6 gives more implementation details of HypDAE. Sec 3.2 & Sec 4.1
Sec. 7 provides the mathematical formulae used in hyperbolic neural networks. Sec 3.2 & Sec 3.3
Sec. 8 shows more results of the ablation study of HypDAE. Sec 4.3
Sec. 9 shows more comparison between the latent manipulation in hyperbolic and Euclidean space. Sec 4.3
Sec. 10 provides examples to show the exceptional out-of-distribution few-shot image generation ability. Sec 4.4
Sec. 11 shows the images generated with different radii in the Poincaré disk. Sec 4.3
Sec. 12 compares the images generated by state-of-the-art few-shot image generation method, i.e. WaveGAN [71], HAE [41] and our methods HypDAE. Sec 4.4
Sec. 13 gives more details of the user study we conducted. Sec 4.4
Sec. 14 gives more examples generated by HypDAE. Sec 4.4
@@ 6 Implementation Details and Analysis
Stage I. As mentioned in Sec 3.2, this stage does not require class labels for the images. To promote diversity, we use only the CLIP image encoder’s class token (dimension $1\times 1024$) for a compact representation, aligning it with the CLIP text feature space via a single-head 5-layer Transformer block following the same settings in [70] that inject features into the diffusion process through cross-attention to replace the text feature in the original stable diffusion model.
@@
Stage II. This stage is the only stage that requires the class labels for given images to learn the hierarchical representation. Although class labels are required in Stage II, the model only needs a small number of labeled data for pre-training and can be fine-tuned on larger datasets without supervision. Furthermore, we show exceptional out-of-distribution generation ability in Sec. 10. For the hyperbolic encoder mentioned in Sec 3.2, we use a single-head 5-layer Transformer block to reduce the dimensionality of the Euclidean latent vector $\boldsymbol{c}$, which is then mapped to hyperbolic space via an exponential map. A hyperbolic feed-forward layer [20] produces the final hierarchical representation $z_{\mathbb{D}}$:
@@
[…]
@@
Stage III. As mentioned in Sec 3.2, this stage does not require class labels for the images. Recall that in Sec 3.2, in the Hyperbolic Delta Mapper framework, we extract delta image-text features, $\Delta i=i_{2}-i_{1}$, and $\Delta t=t_{2}-t_{1}$, as condition vectors during training and inference. As shown in Fig. 2 in the main paper, training involves obtaining CLIP image space direction $\Delta i=i_{2}-i_{1}$ and corresponding hyperbolic direction $\Delta\boldsymbol{c}{h}=\boldsymbol{c}_{h2}-\boldsymbol{c}_{h1}$. A latent mapper network in hyperbolic space, the Hyperbolic Delta Mapper, then predicts the editing direction based on:
$\Delta\boldsymbol{c}_{h}^{\prime}=\textit{HypDeltaMapper}\left(\boldsymbol{c}_{h1},i_{1},\Delta i\right),$ (12)
where $\boldsymbol{c}_{h1}$ and $i_{1}$ are used as the input of Hyperbolic Delta Mapper to provide specialized information for the source image. The Hyperbolic Delta Mapper is divided into three parts: Hyperbolic Mapper, CLIP Mapper, and Fusion Mapper. All Mappers are designed as single-head 10-layer transformer block.
In Stage III, both the CLIP image encoder and CLIP text encoder are loaded during the training process. Besides, the CLIP encoders are frozen, and only the lightweight Hyperbolic Delta Mapper is trainable. We use 1 $\times$ NVIDIA RTX 4090 (24GB) GPU for training, and the batch size is selected as $32$. The $\lambda$ in Eq. (9) in the main paper is selected as $0.1$. We choose the AdamW [38] optimizer and set the learning rate as $1\mathrm{e}{-3}$. A linear learning rate scheduler is used with a step size equal to $5000$, with a multiplier $\gamma=0.5$. We train about $1\mathrm{e}{5}$ steps to get the model to converge on each dataset.
@@
Although training our model requires considerable computing resources as mentioned before, the runtime cost and resources required for the inference stage are affordable. Our model can be inference on a single NVIDIA RTX 4090 GPU (24GB) thanks to our multi-stage training/inference since one does not need to load all models simultaneously.
@@ 7 Hyperbolic Neural Networks
[…]
@@
where $\oplus_{c}$ denotes the Möbius addition defined in Eq. 13 with fixed sectional curvature of the space, denoted by $c$.
@@
Hierarchical Data Sampling. As illustrated in Fig. 13, as the latent code $z_{ref}$ of the reference image $x_{ref}$ moving from the edge to the center of the Poincaré disk, the control of the identity of the sampled images becomes weaker and weaker. The identity of the generated images becomes more ambiguous. To conduct hierarchical image sampling in hyperbolic space, one can move the latent code $z_{ref}$ of the reference image $x_{ref}$ towards the origin of the Poincaré disk. Then, sampling latent points among the “children” of the rescaled reference images. In practice, the semantic diversity of the generated images can be controlled either by setting different values of $r_{\mathbb{D}}$, then calculating $r_{\mathbb{C}}$ based on $r_{\mathbb{D}}$ or setting different values of $r_{\mathbb{C}}$ directly.
@@
where $\oplus_{c}$ denotes the Möbius addition with aforementioned sectional curvature $c$. Therefore, for hierarchical data sampling, we can first define the value of $r_{\mathbb{C}}$, then sample random data points in hyperbolic space. If the distance between the sampled data point and $z_{ref}$ is equal to or shorter than the value, we accept the data point. Otherwise, we move the latent codes along the geodesic between the sampled data point and $z_{ref}$ until the distance is within the scope we define. We show more hierarchical image sampling examples in Sec. 10.
@@ 8 Ablation Study
There are few hyperparameters of HypDAE that control the generation quality and diversity. We conduct ablation studies on each of them in this section.
Hyperbolic Radius. By varying the radii of “parent” images, HypDAE controls the semantic diversity of generated images (Fig. 22), where the “parent” images can be viewed as the image with the average attributes of its children. The quantitative results of the Flowers dataset are presented in Tab. 4. We can see that the diversity increases as the radius becomes smaller, therefore, the value of LPIPS increases accordingly. However, changing too many attributes changes the identity or category of the given images, therefore, the FID decreases when the radius is smaller than $5.5$. In practice, we select $5.5$ as the radius of the parent images for few-shot image generation.
@@
where $\epsilon_{\mathrm{prd}}$, $\epsilon_{\mathrm{uc}}$, $\epsilon_{\mathrm{c}}$, $s$ are the model’s final output, unconditional output, conditional output, and a user-specified weight respectively. The visualizations are shown in Fig. 14, and quantitative results for the Flowers dataset are presented in Tab. 5. Consistent with findings in Tab. 4, diversity, measured by LPIPS, increases as the cfg scale grows. However, excessive cfg scaling can alter the identity or category of the input images, leading to a decline in FID when the cfg scale exceeds $1.3$ Based on these results, we select a cfg scale of $1.3$ to achieve optimal few-shot image generation with a balance between fidelity and diversity.
@@
While lower encoding strength preserves the identity and style of the input images, it reduces diversity. This trade-off is visualized in Fig. 15, and quantitative results for the Flowers dataset are presented in Tab. 6. Consistent with Tab. 4, diversity, measured by LPIPS, increases with higher encoding strength, while excessive encoding strength can cause changes in identity or category. This is reflected in a decrease in FID when encoding strength exceeds $0.95$ (i.e., $5\%$ of the information is encoded). Based on these findings, we set the encoding strength of the stochastic encoder to $0.95$ to achieve reliable few-shot image generation.
@@
Figure 15: Ablation study on the influence of the encoding strength of the stochastic encoder on FFHQ (strength equals $1$ means $x_{0}$ is fully deconstructed, i.e., $x_{T}$ is a Gaussian noise).
@@ 9 Comparison with Euclidean space
In this section, we present a detailed comparison of different latent spaces, as shown in Fig. 16. Compared to classical Euclidean space, hyperbolic space enables smoother transitions between two given images. In hyperbolic space, identity-irrelevant features transition first, followed by a gradual change in identity-relevant features. In contrast, Euclidean space exhibits simultaneous changes in both identity-relevant and identity-irrelevant features, leading to less structured transitions.
These results confirm that our method effectively learns hierarchical representations in hyperbolic space, enabling few-shot image generation by selectively modifying category-irrelevant features—a capability that Euclidean space cannot achieve. Additional interpolation results, provided in Fig. 17, demonstrate that smooth and distortion-free transitions are achievable in hyperbolic space. These findings highlight that HypDAE enables precise geodesic and hierarchical control during editing, offering a significant advantage over traditional approaches.

### v2

@@ 4.5 Few-Shot Image Generation
As described in Sec.˜3, our approach enables few-shot image generation by varying the stochastic subcode $z_{T}$ or shifting the latent code in a randomly chosen semantic direction within the category cluster. We set $r_{\mathbb{D}}=5.5$ and an encoding strength of $0.95$ (i.e., encoding $5\%$ of the information) to achieve this, as illustrated in Fig.˜4. We conduct three experiments demonstrating that HypDAE achieves promising few-shot generation, with additional examples in the SM.
Hyp Radius 6.2 6.0 5.5 5.0 4.5 4.0 3.0
FID($\downarrow$) 15.18 14.67 14.31 14.56 14.71 16.34 20.65
LPIPS($\uparrow$) 0.7035 0.7278 0.7415 0.7643 0.7935 0.8378 0.8964
Table 2: Ablation study of different radii on Animal Faces.
Hyp Radius 6.2 6.0 5.5 5.0 4.5 4.0 3.0
CLIP-S 77.37 75.41 75.15 72.16 71.45 68.20 67.89
CLIP-P 69.62 69.89 72.35 72.86 74.25 76.34 77.00
@@
Quantitative Comparison with State-of-the-Art. To assess fidelity and diversity, we compute FID [22] and LPIPS [64] following the one-shot settings in [37]. Results in Tab.˜1 show significant improvements in FID and LPIPS scores on all four datasets except for the FID score on vggfaces due to the low resolution, affirming the advantages of our diffusion-based approach over previous GAN-based methods.
Qualitative Evaluation. We compare HypDAE qualitatively with WaveGAN [61] and HAE [37]. As shown in Fig.˜9, HypDAE generates highly diverse images with fine-grained fidelity, such as detailed feather textures, which other methods fail to preserve. A user study further evaluates synthesis quality, identity preservation, and category-irrelevant diversity across these methods. As reported in Tab.˜4, HypDAE consistently outperforms all baselines, with notable improvements in fidelity and diversity, attributed to its effective disentanglement of identity-relevant and irrelevant features. Additional details are provided in Sec. I of the SM.
@@ 5 Conclusion
In this work, we present HypDAE, the first diffusion-based method for editing hierarchical attributes in hyperbolic space. By learning the semantic hierarchy of images, our diffusion autoencoder enables continuous and flexible editing of hierarchical features. Experimental results show that HypDAE significantly outperforms existing GAN-based approaches without human-annotated labels, excelling in both random and customized few-shot generation. Furthermore, it facilitates hierarchical image generation with varying semantic similarity, advancing the state of the art in few-shot image generation. We hope this work inspires future research in this direction.
Acknowledgement. This work is supported by the Shun Hing Institute of Advanced Engineering (SHIAE) Fund (No. 8115074).
@@
Sec.˜6 gives more implementation details of HypDAE. Sec 3.2 & Sec 4.1
Sec.˜7 gives detailed explanation of diffusion models.
Sec.˜8 provides the mathematical formulae used in hyperbolic neural networks. Sec 3.2 & Sec 3.3
Sec.˜9 shows more results of the ablation study of HypDAE. Sec 4.3
Sec.˜10 shows more comparisons between the latent manipulation in hyperbolic and Euclidean space. Sec 4.3
Sec.˜11 provides examples to show the exceptional out-of-distribution few-shot image generation ability. Sec 4.4
Sec.˜12 shows the images generated with different radii in the Poincaré disk. Sec 4.3
Sec.˜13 compares the images generated by state-of-the-art few-shot image generation method, i.e. WaveGAN [61], HAE [37] and our methods HypDAE. Sec 4.4
Sec.˜14 gives more details of the user study we conducted. Sec 4.4
Sec.˜15 gives more examples generated by HypDAE. Sec 4.4
@@ 6 Implementation Details and Analysis
Stage I. As mentioned in Sec 3.2, this stage does not require class labels for the images. To promote diversity, we use only the CLIP image encoder’s class token (dimension $1\times 1024$) for a compact representation, aligning it with the CLIP text feature space via a 5-layer fully connected MLP following the same settings in [60] that inject features into the diffusion process through cross-attention to replace the text feature in the original stable diffusion model.
@@
Stage II. This stage is the only stage that requires the class labels for given images to learn the hierarchical representation. Although class labels are required in Stage II, the model only needs a small number of labeled data for pre-training and pseudo labels can be predicted by CLIP as shown in Sec.˜6. Furthermore, we show exceptional out-of-distribution generation ability in Sec.˜11. For the hyperbolic encoder mentioned in Sec 3.2, we use a single-head 5-layer Transformer block to reduce the dimensionality of the Euclidean latent vector $\boldsymbol{c}$ from $1\times 1024$ to $1\times 512$, which is then mapped to hyperbolic space via an exponential map. A hyperbolic feed-forward layer [16] produces the final hierarchical representation $z_{\mathbb{D}}$:
@@
where $\operatorname{E}$ is the Transformer encoder and $f^{\otimes_{c}}$ is the Möbius translation of feed-forward layer $f$ as the map from Euclidean space to hyperbolic space, denoted as Möbius linear layer. In order to perform multi-class classification on the Poincaré disk defined in Sec 3.1, one needs to generalize multinomial logistic regression (MLR) to the Poincaré disk defined in [16]. An extra linear layer needs to be trained for the classification, and the details on how to compute softmax probability in hyperbolic space are shown in Sec.˜8. As mentioned in Sec 3.1, the distance between points grows exponentially with their radius in the Poincaré disk. In order to minimize Eq. (5) in the main paper, the latent codes of fine-grained images will be pushed to the edge of the ball to maximize the distances between different categories while the embedding of abstract images (images have common features from many categories) will be located near the center of the ball. Since hyperbolic space is continuous and differentiable, we are able to optimize Eq. (5) with stochastic gradient descent, which learns the hierarchy of the images.
@@
Although training our model requires considerable computing resources as mentioned before, the runtime cost and resources required for the inference stage are affordable. Our model can inference on a single NVIDIA RTX 4090 GPU (24GB) thanks to our multi-stage training/inference since one does not need to load all models simultaneously.
Pseudo-Labeling. For Flowers, Animal Faces and NABirds, we utilize the CLIP ViT-B/32 model. Given an image ($x$), we extract its embedding using the CLIP image encoder. Similarly, we compute embeddings for a predefined set of class names ($y_{i}$) using the CLIP text encoder. The cosine similarity between the image embedding and each class embedding is computed as:
$\operatorname{sim}(x,y_{i})=\frac{f(x)\cdot g(y_{i})}{|f(x)||g(y_{i})|},$ (10)
where ($f(\cdot)$) and ($g(\cdot)$) denote the CLIP image and text encoders, respectively. The softmax function is applied to convert similarity scores into probabilities. The pseudo-label ($y$) is assigned as the class with the highest probability:
$y=\underset{j}{\mathrm{argmax}}\,\frac{\exp(\operatorname{sim}(x,y_{i}))}{\sum_{j}\exp(\operatorname{sim}(x,y_{j}))}.$ (11)
For the VGGFaces dataset, we employ the DeepFace framework with the VGG-Face architecture. DeepFace predicts pseudo-labels by comparing the face embedding of an image with embeddings of known identities in the dataset. Specifically, we construct a reference database by selecting one image per class, and each input image is assigned to the class with the highest similarity score. This approach ensures robust pseudo-labeling by leveraging DeepFace’s face recognition capabilities.
Dataset Settings We evaluate our method on Animal Faces [39], Flowers [43], VGGFaces [45], FFHQ [30], and NABirds [58] following the settings described in [14, 37].
7 Additional Background - Diffusion Models
Diffusion Denoising Probabilistic Models (DDPM) [24] are generative latent variable models that aim to model a distribution $p_{\theta}(x_{0})$ that approximates the data distribution $q(x_{0})$ and easy to sample from. DDPMs model a “forward process” in the space of $x_{0}$ from data to noise. This is called “forward” due to its procedure progressing from $x_{0}$ to $x_{T}$. Note that this process is a Markov chain starting from $x_{0}$, where we gradually add noise to the data to generate the latent variables $x_{1},\ldots,x_{T}\in X$. The sequence of latent variables, therefore, follows $q(x_{1},\ldots,x_{t}\mid x_{0})=\prod_{i=1}^{t}q(x_{t}\mid x_{t-1})$, where a step in the forward process is defined as a Gaussian transition $q(x_{t}\mid x_{t-1}):=N(x_{t};\sqrt{1-\beta_{t}}x_{t-1},\beta_{t}I)$ parameterized by a schedule $\beta_{0},\ldots,\beta_{T}\in(0,1)$. When $T$ is large enough, the last noise vector $x_{T}$ nearly follows an isotropic Gaussian distribution.
An interesting property of the forward process is that one can express the latent variable $x_{t}$ directly as the following linear combination of noise and $x_{0}$ without sampling intermediate latent vectors:
$x_{t}=\sqrt{\alpha_{t}}x_{0}+\sqrt{1-\alpha_{t}}w,~~w\sim N(0,I),$ (12)
where $\alpha_{t}:=\prod_{i=1}^{t}(1-\beta_{i})$.
To sample from the distribution $q(x_{0})$, we define the dual “reverse process” $p(x_{t-1}\mid x_{t})$ from isotropic Gaussian noise $x_{T}$ to data by sampling the posteriors $q(x_{t-1}\mid x_{t})$. Since the intractable reverse process $q(x_{t-1}\mid x_{t})$ depends on the unknown data distribution $q(x_{0})$, we approximate it with a parameterized Gaussian transition network $p_{\theta}(x_{t-1}\mid x_{t}):=N(x_{t-1}\mid\mu_{\theta}(x_{t},t),\Sigma_{\theta}(x_{t},t))$. The $\mu_{\theta}(x_{t},t)$ can be replaced [24] by predicting the noise $\epsilon_{\theta}(x_{t},t)$ added to $x_{0}$ using equation 12.
@@ 8 Hyperbolic Neural Networks
where $\|\cdot\|$ denotes the $2$-norm of the vector, and $\langle\cdot,\cdot\rangle$ denotes the Euclidean inner product of the vectors.
@@
where $\oplus_{c}$ denotes the Möbius addition defined in Eq.˜13 with fixed sectional curvature of the space, denoted by $c$.
@@
Hierarchical Data Sampling. As illustrated in Fig.˜14, as the latent code $z_{ref}$ of the reference image $x_{ref}$ moves from the edge to the center of the Poincaré disk, the control of the identity of the sampled images becomes weaker and weaker. The identity of the generated images becomes more ambiguous. To conduct hierarchical image sampling in hyperbolic space, one can move the latent code $z_{ref}$ of the reference image $x_{ref}$ towards the origin of the Poincaré disk. Then, sampling latent points among the “children” of the rescaled reference images. In practice, the semantic diversity of the generated images can be controlled either by setting different values of $r_{\mathbb{D}}$, then calculating $r_{\mathbb{C}}$ based on $r_{\mathbb{D}}$, or by setting different values of $r_{\mathbb{C}}$ directly.
@@
where $\oplus_{c}$ denotes the Möbius addition with aforementioned sectional curvature $c$. Therefore, for hierarchical data sampling, we can first define the value of $r_{\mathbb{C}}$, then sample random data points in hyperbolic space. If the distance between the sampled data point and $z_{ref}$ is equal to or shorter than the value, we accept the data point. Otherwise, we move the latent codes along the geodesic between the sampled data point and $z_{ref}$ until the distance is within the scope we define. We show more hierarchical image sampling examples in Sec.˜11.
@@ 9 Ablation Study
There are a few hyperparameters of HypDAE that control the generation quality and diversity. We conduct ablation studies on each of them in this section.
Hyperbolic Radius. By varying the radii of “parent” images, HypDAE controls the semantic diversity of generated images (Fig.˜24), where the “parent” images can be viewed as the image with the average attributes of its children. The quantitative results of the Flowers dataset are presented in Tab.˜5. We can see that the diversity increases as the radius becomes smaller, therefore, the value of LPIPS increases accordingly. However, changing too many attributes changes the identity or category of the given images, therefore, the FID decreases when the radius is smaller than $5.5$. In practice, we select $5.5$ as the radius of the parent images for few-shot image generation.
@@
where $\epsilon_{\mathrm{prd}}$, $\epsilon_{\mathrm{uc}}$, $\epsilon_{\mathrm{c}}$, $s$ are the model’s final output, unconditional output, conditional output, and a user-specified weight, respectively. The visualizations are shown in Fig.˜16, and quantitative results for the Flowers dataset are presented in Tab.˜6. Consistent with findings in Tab.˜5, diversity, measured by LPIPS, increases as the cfg scale grows. However, excessive cfg scaling can alter the identity or category of the input images, leading to a decline in FID when the cfg scale exceeds $1.3$. Based on these results, we select a cfg scale of $1.3$ to achieve optimal few-shot image generation with a balance between fidelity and diversity.
@@
While lower encoding strength preserves the identity and style of the input images, it reduces diversity. This trade-off is visualized in Fig.˜17, and quantitative results for the Flowers dataset are presented in Tab.˜7. Consistent with Tab.˜5, diversity, measured by LPIPS, increases with higher encoding strength, while excessive encoding strength can cause changes in identity or category. This is reflected in a decrease in FID when encoding strength exceeds $0.95$ (i.e., $5\%$ of the information is encoded). Based on these findings, we set the encoding strength of the stochastic encoder to $0.95$ to achieve reliable few-shot image generation.
Hyperparameter Ablation. To validate the robustness of the trade-off parameter in Eq. (7), we rewrite Eq. (7) as: $\mathcal{L}=\lambda\cdot\mathcal{L}_{\text{hyper}}+\mathcal{L}_{\text{rec}}$, ablate the loss to analyze this trade-off. As shown in Fig.˜15, increasing $\lambda$ improves semantic consistency (lower FID) while reducing diversity (lower LPIPS), validating the controllability introduced by the hyperbolic component.
@@
Figure 15: Ablation Study of trade-off adaptive hyperparameter $\lambda$ in the loss function.
@@
Figure 17: Ablation study on the influence of the encoding strength of the stochastic encoder on FFHQ (strength equals 11 means $x_{0}$ is fully deconstructed, i.e., $x_{T}$ is a Gaussian noise).
@@ 10 Comparison with Euclidean space
In this section, we present a detailed comparison of different latent spaces, as shown in Fig.˜18. Compared to classical Euclidean space, hyperbolic space enables smoother transitions between two given images. In hyperbolic space, identity-irrelevant features transition first, followed by a gradual change in identity-relevant features. In contrast, Euclidean space exhibits simultaneous changes in both identity-relevant and identity-irrelevant features, leading to less structured transitions.
These results confirm that our method effectively learns hierarchical representations in hyperbolic space, enabling few-shot image generation by selectively modifying category-irrelevant features—a capability that Euclidean space cannot achieve. Additional interpolation results, provided in Fig.˜19, demonstrate that smooth and distortion-free transitions are achievable in hyperbolic space. These findings highlight that HypDAE enables precise geodesic and hierarchical control during editing, offering a significant advantage over traditional approaches.